OpenStudy (anonymous):

Chuck's CheeseHouse makes two types of cheese; mozzarella and parmesan. Mozzarella costs $2.60 per pound and parmesan costs $3.30 per pound. How much parmesan cheese went into making a 60 pound mixture worth $3.10 per pound? Round to the nearest pound. 17 lbs 23 lbs 43 lbs 37 lbs

OpenStudy (ganpat):

let mozzarella = x & parmesan = y so 2.6x + 3.3y = 60(3.1) and x + y = 60.. can u solve them ??

OpenStudy (anonymous):

??? still nothing

OpenStudy (ganpat):

two equations.. solve them simultaneously.. x + y = 60.. ----> x = 60 - y substitute in equation.. 2.6x + 3.3y = 60(3.1) 2.6 (60 - y) + 3.3y = 186.. 156 - 2.6y + 3.3y = 186 3.3y - = 186 - 156 0.7y = 30 so y = 42.8 approx = 43.. thats ur answer...